From a group of 5 men and 4 women, a comittee of 6 is to be selected. How many different selections can be formed having exactly 4 men and 2 women?
Answer by stanbon(75887)

From a group of 5 men and 4 women, a comittee of 6 is to be selected. How many different selections can be formed having exactly 4 men and 2 women?
----
Ans: 5C4*4C2 = 5*6 = 30
===============================
Cheers,
Stan H.